You are buying an automobile that costs $15,000. You are paying $5,000 immediately and the remaining $10,000 in four annual eYou are buying an automobile that costs $15,000. You are paying $5,000 immediately   and the remaining $10,000 in four annual end-of-year principal payments of $2,500 each. In addition to the $2,500, you must pay 12% interest on the unpaid balance of the loan each year. Prepare a cash flow table to represent your cash outlay per year.
Answer #1

In year 0 you will pay $ 5,000 and in other years you will pay a principal amount of $ 2,500 each year along with interest of 12% on unpaid balance.

In year 1, unpaid balance is $ 10,000 interest on it will be $ 10,000 × 0.12 = $ 1,200

Payment of 1st year = 2,500 + 1,200 = $ 3,700

In year 2, loan balance = $ 10,000 - 2,500 = $ 7,500

Interest of second year = 0.12 × 7,500 = $ 900

Payment of 2nd year = $ 2,500 + 900 = $ 3,400

Like wise we can calculate for the remaining years.

Refer the attached table for cash outlay each year

1 Year BOY Principal Interest Cash Outflow 0 $ 5,000 | $ 5,000 $ 10,000 $ 2,500 $ 1,200 $ 3,700 2 $ 7,500 $ 2,500 $ 900 $ 3,4

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.